please help me with this equation
b-(1-2b)+(b-3)=-4

Answer by gonzo(654)

b - (1 - 2*b) + (b - 3) = -4
becomes
b - 1 + 2*b + b - 3 = -4
becomes
4*b -4 = -4
becomes
4*b = 0
becomes
b = 0
-------------------------------
if you minus a minus you get a plus